# End-to-End Large Portfolio Optimization for Variance Minimization with Neural Networks through Covariance Cleaning

### v2

@@ 4.2 Training Procedure
The proposed architecture is designed to be independent of the number of assets $n$. This means that the same model can be applied to different subsets of assets without the need for retraining. However, we know from RMT that the filtering should depend on the aspect ratio $q=n/\Delta t_{\text{in}}$; therefore, during the training phase, the NN should be fed with examples of different $n$ sizes if we aim to obtain a model that generalizes well to larger portfolio sizes. This is done by randomly and uniformly sampling the number of assets $n$ from a [50,350][50,350][ 50 , 350 ] range in every batch of the training. The number of days $\Delta t_{\text{in}}$ is fixed to 120012001200 trading days; in fact, the presence of the module of Sec. 3.1 includes $2\Delta t_{\text{in}}$ trainable variables $(\alpha_{t},\beta_{t})$, which are associated with specific past lags. In the future version, this could be generalized to a variable number of past lags, but it would require a more complex architecture.
@@
We trained our model on a dataset of US equities spanning January 1990 to December 2024. For each validation date $t$, we define the calibration window as the period from the dataset start through December 31 of the year preceding $t$. In particular, we conduct validation over the interval [2000-01-01, 2024-12-31], resulting in 24 distinct trained NNs. To construct a robust training procedure, we randomly draw a date $t$ such that the interval $[t-\Delta t_{\mathrm{in}}-1,t+\Delta t_{\mathrm{out}}]$ is fully contained within the calibration window. We then randomly select $n$ assets from the investable universe at $t-1$ (see Section 4.1). The NN is trained on the adjusted close-to-close returns over the input window $[t-\Delta t_{\mathrm{in}},t-1]$ together with the realized OOS covariance matrix $\bm{\Sigma}_{\mathrm{out}}$, computed from the adjusted close-to-close returns over $[t,\,t+\Delta t_{\mathrm{out}}]$. A one-day shift between the training inputs and the OOS returns is introduced to avoid data leakage, since the return at $t$ depends on the closing price at the end of day $t$, which is not available intra-day.
@@
According to the training procedure, the probability that NN will be trained on the same dataset is very low for $n$ sufficiently smaller than the investable universe on a given day. Consequently, continuously monitoring validation-set performance for early stopping becomes less critical. However, to further reduce the risk of overfitting, we used a decay factor of the learning rate, which is set to $10^{-4}$ and drops every batch by a factor of $0.99^{1/500}$, and we clip the gradients by norm to 1.01.01.0. The gradient is then optimized with the Adam optimizer with an automatic rescaling of the loss function. We trained all our models with 100100100 epochs, with 500500500 steps per epoch, and a batch size of 323232 samples. To ensure that the convergence is stable, for each of the 24 years, we trained our NN 10 times independently, so finally, we obtained 240 trained NNs.
@@
[…]
@@ 4.3.1 Lag-Transformed Returns
After the calibration, we can analyze the learned weighting factors $\alpha_{t}$ and $\beta_{t}$ of Eq. (14) as a function of the lag $t$ to understand how the NN is weighting and transforming the past returns. Given that we ran 101010 independent training runs for each of the 242424 years of the validations, we obtained 240240240 trained models. We can then compute the average of the weighting factors over the trained models, which are shown in Fig. 3.
@@
Figure 3: The upper plots show the calibrated weighting factors $\alpha_{t}$ and $\beta_{t}$ as a function of the lag. The factors shown here are an average over the 240 trained models. Low values of the $x$-axis correspond to the most recent past, while high values correspond to the oldest past. Both figures show an attempt to fit them with a power law for $\alpha_{t}$ and with a saturating exponential for $\beta_{t}$ indicated with the dashed red lines. The lower plots show an example of the return transformation of the annualized returns for the last trained model over past examples. The left plot refers to the day before the investment, and the right plot refers to the first day in the dataset. The marginal distribution shows how the empirical distribution is transformed.
[…]
@@
The lower panels of Fig. 3 present the combined impact of the shrinkage and sigmoid transformations. In the remote past, the transformed values not only contract toward zero but also assume effectively binary states. As a result, beyond approximately one year of lag the series conveys only the sign of daily returns. When the Pearson correlation is applied to a binarized series, it coincides with the Phi coefficient [68, 69]. In contrast, the most recent returns still exhibit a strong binary character yet develop a pronounced uniform region around the center. This feature reflects the fact that the sigmoid kernel closely tracks the empirical cumulative distribution. By transforming the raw data via the empirical CDF, the Pearson correlation of the transformed series becomes the Spearman coefficient [70]. However, this exact equivalence holds only when each return series is rank-transformed by its own empirical CDF, whereas using a single merged CDF across all assets only approximates the Spearman coefficient and remains inexact due to volatility-driven distortions across assets. In summary, as the lag increases, the network’s operative correlation metric shifts continuously from a Spearman-type measure at short lags toward a Phi-type measure at long lags.
@@ 4.3.2 Eigenvalue Sensitivity Analysis
[…]
In the left panel of Fig. 4, we present the eigenvalue spectra for the MLE, the QIS shrinkage, and the NN-transformed eigenvalues $\bm{\lambda}_{\text{NN}}$ after inversion of Eq. (17). The range of the largest eigenvalues coincides across all three estimators, whereas the bulk of the spectrum exhibits the most pronounced differences. The MLE assigns the lowest values to the bulk eigenvalues. The QIS shrinkage attenuates this dispersion by smoothing the spectrum. Neither method reproduces the abrupt truncation characteristic of eigenvalue clipping [5]; instead, while both the MLE and QIS yield a bulk spectrum spanning approximately one order of magnitude, the NN transformation compresses the bulk eigenvalues to within roughly half an order of magnitude. This compression causes the NN transformation to resemble more the abrupt truncation of eigenvalue clipping.
The right panel of Fig. 4 shows the standard deviation of the log-transformed eigenvalues across calibration samples. All estimators exhibit pronounced variability for the largest eigenvalues. In the bulk of the spectrum, the MLE and the QIS continue to display substantial dispersion, whereas the NN-transformed eigenvalues form a nearly flat plateau. This plateau indicates that the NN mapping renders the bulk eigenvalues effectively insensitive to the input data, akin to the asset-agnostic behavior of the AO estimator, while still responding to the extreme eigenvalues.
@@ 4.4 Backtesting Procedure
The calibrated network is trained to predict the GMV portfolio by allowing unrestricted short positions. This design choice yields a closed-form mapping of the optimal weights via the analytical expression in equation Eq. (1). If one imposes a non-negativity constraint on the weights, the GMV solution must be obtained through a numerical solver. Although specialized NN layers can implement such constrained quadratic programs with end-to-end differentiation [71], this approach incurs substantially higher computational cost and, in our tests, suffers from vanishing gradients. On the other side, allowing short sales solely to minimize variance might be undesirable, since it amplifies gross leverage and incurs additional borrowing and transaction costs.
To reconcile analytical tractability with realistic trading constraints, after the uncontrained calibration, we can alter the architecture so that the network directly outputs an estimate of the inverse covariance matrix, $\bm{\Sigma}_{\mathrm{NN}}^{-1}$ (see Fig. 1). We invert this matrix and then compute portfolio weights under a long-only constraint by solving the QP externally.
In Sec. 4.4.2 we evaluate our method against alternative estimators defined in Sec. 4.4.1 under idealized, frictionless conditions. In this experiment, we calibrate the network without trading constraints, i.e., allowing both long and short positions. Then, we test both the raw uncontrained output and the long-only version described above. In Sec. 4.4.3 investigates the model’s extrapolation performance. We train on a small universe of assets and then test on a substantially larger universe, enforcing long-only positions and employing an operational simulator that incorporates realistic transaction costs.
@@ 4.4.1 Benchmarking Alternative Estimators
We compare two broad classes of portfolio estimators, (i) univariate benchmarks and (ii) multivariate covariance-cleaning methods, across two complementary experimental designs introduced in Secs.  4.4.2 and 4.4.3. In Sec.  4.4.2 conduct a bootstrap-style backtest without transaction costs to enable frictionless, synthetic comparisons over varying universe sizes. In Sec. 4.4.3, we perform a single, realistic simulation on the top-capitalisation universe of 1,000 stocks, incorporating trading frictions, liquidity constraints, and slippage. Univariate benchmarks rely solely on individual asset variances or market–weight information. Multivariate methods instead apply covariance-cleaning procedures to the estimated correlation matrix prior to optimization. Although univariate strategies are pervasive in index construction and passive instruments, their ex post performance is predictably inferior; accordingly, we include them only in our detailed analysis of the realistic backtest of Sec. 4.4.3.
@@
Multivariate benchmarks differ only in the way the $n\times n$ correlation matrix $\widehat{\bm{C}}=\bm{C}_{\bullet}$ is obtained. After the correlation matrix is estimated, we always project on the univariate volatilities $\widehat{\bm{\Sigma}}=\bm{D}_{\textrm{MLE}}\bm{C}_{\bullet}\bm{D}_{\textrm{MLE}}$ and we apply the QP optimization. Filtering the correlation rather than the covariance is a necessary step to put RIEs in their best performance setup. The simplest method is the sample MLE uses the unfiltered sample correlation matrix computed over the look-back window. The Power Mapping (PM), introduced by Guhr and Kälber [74], suppresses estimation noise by raising every off-diagonal element of the de-meaned correlation matrix to a power $\gamma>1$ and re-normalising. For the LS, we use the standard implementation of Sklearn [75] where the parameter is set up according to the original work of Ledoit and Wolf [6]. For QIS, we used a Python implementation released in the GitHub repository [76] that we adapted to improve the computational performance. Finally, AO is calibrated using the universe of stocks as defined in Sec. 4.1 within 1990-2000. Then the eigenvalues are kept constant for the whole validation period, 2000-2024, independently of the stock in the basket.
We exclude dynamical models, most notably the DCC estimator and its myriad extensions, from our benchmark analysis. Although a vast plethora of DCC variants has emerged in the literature (see Ref. [77]), a thorough comparison against all these methods would require a dedicated study: the differing model formulations and hyper-parameter protocols alone would demand a tailored experimental framework, and the well-documented computational overhead and convergence instabilities of DCC algorithms further compound the effort. More fundamentally, none of the estimators considered in this work, including our end-to-end NN architecture, are designed to produce genuinely time-evolving covariance matrices; they all operate within a static eigenvalue-filtering and shrinkage paradigm. Incorporating DCC-based methods would therefore conflict with the conceptual scope of our paper. A coherent evaluation would entail extending our network with modules to map sequences of covariance matrices and then re-benchmarking it against the full spectrum of DCC approaches, a task beyond the present work. Moreover, Ref. [27] already shows that a standard DCC implementation underperforms the AO estimator on the same universe of the top 1,000 U.S. large-cap equities, with calibration window $\Delta t_{\textrm{in}}=1200$, yielding consistently lower Sharpe ratios for DCC in realistic settings. This evidence further justifies our focus on static cleaning and shrinkage procedures.

### v3

@@ 4.2 Training Procedure
The proposed architecture is designed to be independent of the number of assets $n$. This means that the same model can be applied to different subsets of assets without the need for retraining. However, we know from RMT that the filtering should depend on the aspect ratio $q_{\textrm{in}}:=n/\Delta t_{\text{in}}$; therefore, during the training phase, the NN should be fed with examples of different $n$ sizes if we aim to obtain a model that generalizes well to larger portfolio sizes. Accordingly, at each batch we randomly and uniformly sampling the number of assets $n$ from a $[50,350]$ range in every batch of the training. This resampling also enlarges the effective training set, because the number of possible cross sections grows combinatorially and repeating the exact same asset combination becomes unlikely when the investable universe is large. The number of days $\Delta t_{\text{in}}$ is fixed to $1200$ trading days; in fact, the presence of the module of Sec. 3.1 includes $2\Delta t_{\text{in}}$ trainable variables $(\alpha_{t},\beta_{t})$, which are associated with specific past lags. In the future version, this could be generalized to a variable number of past lags, but it would require a more complex architecture.
@@
We trained our model on a dataset of US equities spanning January 1990 to December 2024. For each validation date $t$, we define the calibration window as the period from the dataset start through December 31 of the year preceding $t$. In particular, we conduct validation over the interval [2000-01-01, 2024-12-31], resulting in 24 distinct trained NNs. To construct a robust training procedure, we randomly draw a date $t$ such that the interval $t-\Delta t_{\mathrm{in}}-1,t+\Delta t_{\mathrm{out}}]$ is fully contained within the calibration window. We then randomly select $n$ assets from the investable universe at t1t-1 (see Section 4.1). The NN is trained on the adjusted close-to-close returns over the input window $t-\Delta t_{\mathrm{in}},t-1]$ together with the realized OOS covariance matrix $\bm{\Sigma}_{\mathrm{out}}$, computed from the adjusted close-to-close returns over $t,\,t+\Delta t_{\mathrm{out}}]$. A one-day shift between the training inputs and the OOS returns is introduced to avoid data leakage, since the return at $t$ depends on the closing price at the end of day $t$, which is not available intra-day.
@@
According to the training procedure, the probability that NN will be trained on the same dataset is very low for $n$ sufficiently smaller than the investable universe on a given day. Consequently, continuously monitoring validation-set performance for early stopping becomes less critical. However, to further reduce the risk of overfitting, we used a decay factor of the learning rate, which is set to $10^{-4}$ and drops every batch by a factor of $0.99^{1/500}$, and we clip the gradients by norm to $1.0$. The gradient is then optimized with the Adam optimizer with an automatic rescaling of the loss function. We trained all our models with $100$ epochs, with $500$ steps per epoch, and a batch size of $32$ samples. To ensure that the convergence is stable, for each of the 24 years, we trained our NN 10 times independently, so finally, we obtained 240 trained NNs.
All training runs were executed on CPU, without GPU acceleration, and did not pose any computational bottlenecks. On a standard Intel i7 processor, one epoch (500 steps with batch size 32) requires approximately 200 seconds on average. After calibration, inference is comparatively lightweight, computing portfolio weights for a single instance requires on the order of $10^{-1}$ seconds on the same hardware.
@@
For further evidence of robustness against overfitting, Fig. 2 displays the evolution of training and validation losses over successive epochs. The loss trajectories remain consistent across independent runs on the same data interval. Both training and validation losses decrease monotonically. A monotonic decline of the training loss is expected; the validation loss, by contrast, often increases once overfitting sets in. After $60$ epochs, the training loss plateaus, exhibiting an average change of less than $3.5\%$ over the final forty epochs. The validation loss continues to decline at a faster pace, falling by approximately $6\%$ during the same period. These results indicate that terminating training around $100$ epochs would probably incur minimal overfitting risk.
@@ 4.3.1 Lag-Transformed Returns
After the calibration, we can analyze the learned weighting factors $\alpha_{t}$ and $\beta_{t}$ of Eq. (14) as a function of the lag $t$ to understand how the NN is weighting and transforming the past returns. Given that we ran $10$ independent training runs for each of the $24$ years of the validations, we obtained $240$ trained models. We can then compute the average of the weighting factors over the trained models, which are shown in Fig. 3.
@@
Figure 3: The upper plots show the calibrated weighting factors $\alpha_{t}$ and $\beta_{t}$ as a function of the lag. The factors shown here are an average over the 240 trained models. Low values of the $x$-axis correspond to the most recent past, while high values correspond to the oldest past. Both figures show an attempt to fit them with a power law for $\alpha_{t}$ and with a saturating exponential for $\beta_{t}$ indicated with the dashed red lines. Min-Max ranges highlight the consistency of the results over the 240 independent training. The lower plots show an example of the return transformation of the annualized returns for the last trained model over past examples. The left plot refers to the day before the investment, and the right plot refers to the first day in the dataset. The marginal distribution shows how the empirical distribution is transformed.
Interestingly, the rescaling factor $\alpha_{t}$ seems to follow a power law with an exponent of approximately 0.2-0.2, which in substance means that the most recent past is weighted slightly more than the oldest past. In fact, half of the contribution is given from the $40\%$ most recent past. The emergence of the power law is quite surprising, being the most widely used decay factor based on Exponentially Weighted Moving Average (EWMA) [38, 71], whereas the power law decay would instead imply a Hyperbolic Weighted Moving Average (HWMA). A model based on HWMA was proposed in Ref. [41], but to our knowledge, it is not very popular yet.
The clipping factor $\beta_{t}$ shows values within the range $[0.4,1.0]$. Interestingly, the clipping factor $\beta_{t}$ seems to saturate for larger lags. To have an idea about the range of clipping, for the farthest data points, assuming $\alpha_{t}=1$, a daily return of $1\%$ is clipped to approximately $0.4\%$ while the same return in the most recent past is clipped to approximately $0.9\%$.
@@
The lower panels of Fig. 3 present the combined impact of the shrinkage and sigmoid transformations. In the remote past, the transformed values not only contract toward zero but also assume effectively binary states. As a result, beyond approximately one year of lag the series conveys only the sign of daily returns. When the Pearson correlation is applied to a binarized series, it coincides with the Phi coefficient [1, 81]. In contrast, the most recent returns still exhibit a strong binary character yet develop a pronounced uniform region around the center. This feature reflects the fact that the sigmoid kernel closely tracks the empirical cumulative distribution. By transforming the raw data via the empirical CDF, the Pearson correlation of the transformed series becomes the Spearman coefficient [24]. However, this exact equivalence holds only when each return series is rank-transformed by its own empirical CDF, whereas using a single merged CDF across all assets only approximates the Spearman coefficient and remains inexact due to volatility‐driven distortions across assets. In summary, as the lag increases, the network’s operative correlation metric shifts continuously from a Spearman‐type measure at short lags toward a Phi‐type measure at long lags.
The above analysis also clarifies an architectural constraint that affects the choice of the IS window length. In the current implementation, the lag-transformation block assigns a distinct pair of parameters $(\alpha_{t},\beta_{t})$ to each look-back lag $t\in\{1,\dots,\Delta t_{\mathrm{in}}\}$ (Eq. (14)), so the number of trainable variables in this module scales as $2\Delta t_{\mathrm{in}}$. Consequently, $\Delta t_{\mathrm{in}}$ must be fixed at training time and is kept unchanged at inference. At the same time, the learned shapes in Fig. 3 suggest a natural route toward flexible window lengths, since $\alpha_{t}$ and $\beta_{t}$ appear to follow smooth, low-complexity patterns (approximately a power law and a saturating exponential, respectively), one could replace the per-lag free parameters by a low-dimensional parametric kernel. Such a formulation would decouple the number of learnable parameters from $\Delta t_{\mathrm{in}}$ and would allow a variable look-back horizons within the same architectural family.
@@ 4.3.2 Eigenvalue Sensitivity Analysis
Figure 4: Eigenvalue sensitivity analysis. Left: median of the eigenvalues as a function of the rank, the colored bands indicate a $95\%$ percentile coverage. Right: Standard deviation of the log eigenvalues as a function of the rank. The eigenvalues are computed on 1024 samples from the full calibration dataset. Abbreviations: NN = Neural Network estimator; MLE = Maximum-Likelihood (sample) Estimator; QIS = Quadratic-Inverse Shrinkage.
In the left panel of Fig. 4, we present the eigenvalue spectra for the MLE, the QIS shrinkage, and the NN‐transformed eigenvalues $\bm{\lambda}_{\text{NN}}$ after inversion of Eq. (17). The range of the largest eigenvalues coincides across all three estimators, whereas the bulk of the spectrum exhibits the most pronounced differences. The MLE assigns the lowest values to the bulk eigenvalues. The QIS shrinkage attenuates this dispersion by smoothing the spectrum. Neither method reproduces the abrupt truncation characteristic of eigenvalue clipping [43]; instead, while both the MLE and QIS yield a bulk spectrum spanning approximately one order of magnitude, the NN transformation compresses the bulk eigenvalues to within roughly half an order of magnitude. This compression causes the NN transformation to resemble more the abrupt truncation of eigenvalue clipping.
The right panel of Fig. 4 shows the standard deviation of the log‐transformed eigenvalues across calibration samples. All estimators exhibit pronounced variability for the largest eigenvalues. In the bulk of the spectrum, the MLE and the QIS continue to display substantial dispersion, whereas the NN‐transformed eigenvalues form a nearly flat plateau. This plateau indicates that the NN mapping renders the bulk eigenvalues effectively insensitive to the input data, akin to the asset‐agnostic behavior of the AO estimator, while still responding to the extreme eigenvalues.
@@ 4.4 Backtesting Procedure
The calibrated network is trained to predict the GMV portfolio by allowing unrestricted short positions. This design choice yields a closed‐form mapping of the optimal weights via the analytical expression in equation Eq. (1). If one imposes a non‐negativity constraint on the weights, the GMV solution must be obtained through a numerical solver. Although specialized NN layers can implement such constrained quadratic programs with end‐to‐end differentiation [2], this approach incurs substantially higher computational cost and, in our tests, suffers from vanishing gradients. On the other side, allowing short sales solely to minimize variance might be undesirable, since it amplifies gross leverage and incurs additional borrowing and transaction costs.
To reconcile analytical tractability with realistic trading constraints, after the uncontrained calibration, we can alter the architecture so that the network directly outputs an estimate of the inverse covariance matrix, $\bm{\Sigma}_{\mathrm{NN}}^{-1}$ (see Fig. 1). We invert this matrix and then compute portfolio weights under a long‐only constraint by solving the QP externally.
In Sec. 4.4.2 we evaluate our method against alternative estimators defined in Sec. 4.4.1 under idealized, frictionless conditions. In this experiment, we calibrate the network without trading constraints, i.e., allowing both long and short positions. Then, we test both the raw uncontrained output and the long-only version described above. In Sec. 4.4.3 investigates the model’s extrapolation performance. We train on a small universe of assets and then test on a substantially larger universe, enforcing long‐only positions and employing an operational simulator that incorporates realistic transaction costs.
@@ 4.4.1 Benchmarking Alternative Estimators
We compare two broad classes of portfolio estimators, (i) univariate benchmarks and (ii) multivariate covariance‐cleaning methods, across two complementary experimental designs introduced in Secs.  4.4.2 and 4.4.3. In Sec.  4.4.2 conduct a bootstrap‐style backtest without transaction costs to enable frictionless, synthetic comparisons over varying universe sizes. In Sec. 4.4.3, we perform a single, realistic simulation on the top‐capitalisation universe of 1,000 stocks, incorporating trading frictions, liquidity constraints, and slippage. Univariate benchmarks rely solely on individual asset variances or market–weight information. Multivariate methods instead apply covariance‐cleaning procedures to the estimated correlation matrix prior to optimization. Although univariate strategies are pervasive in index construction and passive instruments, their ex post performance is predictably inferior; accordingly, we include them only in our detailed analysis of the realistic backtest of Sec. 4.4.3.
@@
Multivariate benchmarks differ only in the way the $n\times n$ correlation matrix $\widehat{\bm{C}}=\bm{C}_{\bullet}$ is obtained. After the correlation matrix is estimated, we always project on the univariate volatilities $\widehat{\bm{\Sigma}}=\bm{D}_{\textrm{MLE}}\bm{C}_{\bullet}\bm{D}_{\textrm{MLE}}$ and we apply the QP optimization. Filtering the correlation rather than the covariance is a necessary step to put RIEs in their best performance setup. The simplest method is the sample MLE uses the unfiltered sample correlation matrix computed over the look-back window. The Power Mapping (PM), introduced by Guhr and Kälber [32], suppresses estimation noise by raising every off-diagonal element of the de-meaned correlation matrix to a power $\gamma>1$ and re-normalising.
For the LS, we use the standard implementation of Sklearn [42] where the parameter is set up according to the original work of Ledoit and Wolf [48]. For QIS, we used a Python implementation released in the GitHub repository [54] that we adapted to improve the computational performance. Finally, AO is calibrated using the universe of stocks as defined in Sec. 4.1 within 1990-2000. Then the eigenvalues are kept constant for the whole validation period, 2000-2024, independently of the stock in the basket.
We do not aim to provide an exhaustive benchmark of multivariate GARCH/DCC models and their numerous extensions. The DCC literature has produced a wide spectrum of specifications and calibration protocols (see, e.g., Ref. [70]), and a systematic comparison across this full class would require a dedicated experimental framework, especially given the well-documented computational overhead and potential convergence instabilities of high-dimensional DCC estimation. More importantly, our contribution targets a different design point: our end-to-end NN (and most benchmarks in this section) produces a single, window-based covariance estimate per rebalancing date and operates within a static eigenvalue-filtering and shrinkage paradigm, rather than learning an explicit state-space evolution for the covariance process. A fully coherent like-for-like comparison would therefore entail extending the NN with modules that ingest and propagate sequences of covariance states, followed by re-benchmarking against the broader family of dynamic models.
However, because DCC-type models remain the canonical econometric benchmark for time-varying correlations since Engle (2002) [23], we include a representative set of state-of-the-art high-dimensional DCC variants in our experiments of Sec. 4.4.3. Specifically, beyond the baseline DCC, we consider shrinkage-regularised DCC implementations in the spirit of Engle, Ledoit, and Wolf [21], where the correlation-targeting matrix is itself denoised via nonlinear eigenvalue shrinkage (QuEST/NLS), and we also test a variant in which this shrinkage step is replaced by quadratic-inverse shrinkage (QIS) [53]. These specifications are particularly relevant here because they explicitly couple DCC dynamics with rotation-invariant spectral cleaning of the unconditional correlation target. Finally, we include a factor-augmented extension following De Nard, Ledoit, and Wolf [17], in which an observed Fama-French factor is used to model the dominant market-wide component and DCC dynamics are applied to the residual covariance. This augmentation is designed to stabilise large-universe covariance forecasts by preserving economically meaningful factor directions while still allowing time-variation in the idiosyncratic dependence structure. In related large-universe studies, such hybrid DCC-shrinkage and factor-augmented DCC specifications represent the strongest-performing members of the DCC family, yet they remain challenging in portfolio applications due to instability and turnover effects [10].